OpenStudy (anonymous):

will fan and metal PLZ HELP!!! Cole drew a rectangle that is 3 meters by 10 meters. It has an area of 30 square meters and a perimeter of 26 meters. Which rectangle has dimensions that give the same perimeter as Cole's rectangle, but has a different area? A. 15 meters by 2 meters B. 7 meters by 6 meters C. 5 meters by 6 meters D. 2 meters by 13 meters

OpenStudy (anonymous):

so perimeter of a rectangle is determined by 2w+2h, in this case it is 2h+2w= 26 which of these options fits

OpenStudy (anonymous):

so whats the awnser i cant do this its to hard and this is my last one just need an awnser

OpenStudy (anonymous):

but i am reading and it starts to make sence

OpenStudy (anonymous):

so can you plz help me out @Brill

OpenStudy (anonymous):

Just guess and check, multiply each of the height and width by 2, then add them together and see if they add up to 26

OpenStudy (anonymous):

For example I'll try option A... The height and width are 15 and 2, so 15x2= 30 and 2x2=4.... 30+4=34 So that isnt the right answer

OpenStudy (anonymous):

Just repeat that proccess

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i thinks its c because like my brothers are ting to help

OpenStudy (anonymous):

and they cant

OpenStudy (anonymous):

Ok so let's try what I said to do with C 5x2=10 6x2=12 10+12=22

OpenStudy (anonymous):

So that isnt the right answer

OpenStudy (anonymous):

keep trying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about b

OpenStudy (anonymous):

Yep Because 7x2=14 and 6x2=12 14+12=26

OpenStudy (anonymous):

Good job :D

OpenStudy (anonymous):

YES thank you so much <3